OpenStudy (anonymous):

IF F(x) = x^2/(x^2+4).. A. what function is g(x)? B. What function is f(x)? C. Verify that f(g(x)) = F(x)

OpenStudy (tkhunny):

That's a very poor question. I think it can be decoded like this... g(x) = x^2 f(x) = x/(x+4) The nature of the question motivates me just to give you an answer. There are many other answers. That's just one of them that pops out.

OpenStudy (anonymous):

alright thanks